# Replay Consolidation with Label Propagation for Continual Object Detection

### v1

@@ 3.3 Feature Distillation
However, the YOLOv8 architecture includes a neck in addition to the usual backbone $h_{\omega}$. Therefore, we propose to compute the distillation loss on the intermediate features of the backbone and the neck. Recalling that the YOLOv8 backbone outputs three feature maps at multiple levels $C_{3},C_{4},C_{5}$ and the neck does the same $P_{3},P_{4},P_{5}$ [30, 31], the overall loss can be computed as:
@@
3.4 OCDM
The OCDM method was originally proposed to work in the multi-label image classification setting [32]. Its main contribution is proposing a new selection mechanism that differs from the classic random selection used when constructing the replay memory. Referring to Fig. 4, OCDM addresses the unbalanced label distribution found in replay memory samples, which makes them unrepresentative for many non-frequent labels. This is done by ensuring a more uniform label frequency distribution via a greedy approach that removes samples one by one based on how much they contribute to the distribution. We integrate OCDM’s selection mechanism into RCLPOD to determine which samples to include in the replay memory after it has been enhanced by the Label Propagation technique.
@@ 4.1 Scenarios and Metrics
As in [18] to evaluate the performance of YOLOv8, we report the mean average precision (mAP) at a 0.5 IoU threshold (mAP50) for VOC and the mAP weighted across different IoU from 0.5 to 0.95 (mAP50-95) for COCO.
@@ 4.2 YOLO training details
To evaluate how the CL performances are affected by the model size, in our study we consider two different versions of YOLOv8: YOLOv8n with 3.2M parameters and YOLOv8m with 25.9M parameters. We initialize the backbone parameters with the ones pre-trained on Image-Net, available on [14].
In Table 3 we report the hyper-parameters used for training. In particular, by following [14, 17], we use SGD with Nestorov momentum and weight decay. For both the learning rate and the momentum we have 3 warmup epochs, while just for the learning rate we employ a linear decay scheduling from $10^{-2}$ to $10^{-4}$. In all the experiments we set the number of epochs per task to 100, which is a suitable value for reaching convergence.
@@ 4.3 CL strategies details
We compare the following methods against RCLPOD: Joint training, Fine-tuning, Replay, OCDM, LwF, and Pseudo-Label. In particular, while Replay, LwF, and Pseudo-label were already evaluated in the CLOD setting, this is the first time, to the best of our knowledge, that OCDM approach is implemented and tested in the CLOD setting. Join Training is considered as an upper bound and corresponds to the model performances when trained on the entire dataset with all the classes. Fine-tuning is used as a lower bound; each task is trained sequentially, and no CL technique is employed to avoid catastrophic forgetting.
Regarding the Replay methods, the memory has a small capacity fixed to around 5% of the entire dataset, namely 800 images for PASCAL VOC and 6,000 for COCO. Once the model is trained on task $i$, $n$ samples are randomly selected from the current training set and added to the Replay memory. To prevent exceeding the memory’s total capacity when adding $n$ new samples, $n$ samples are randomly removed from the replay memory. For OCDM, we use the same memory size as all the other replay-based approaches.
@@
Fine-Tuning 36.4 16.4 26.5 2.9 13.6 2.9
Replay 54.5 60.9 56.9 40.0 14.1 8.0
LwF 36.7 3.7 13.5 3.1 8.8 1.9
OCDM 56.5 59.9 54.7 50.2 16.7 11.0
Pseudo-label 71.4 46.6 59.7 16.3 24.3 16.1
@@ 5 Results
In this section, we discuss the outcomes of each method on the VOC and COCO CLOD benchmarks. Table 1 reports the results after all tasks have been completed for YOLOv8n.
In Sec. 5.1, we discuss the results obtained in the 2-tasks scenarios, namely 15p5, 10p10, 19p1, and 40p40. In Sec. 5.2, we consider the more challenging scenarios 15p1 and 40p10.
@@ 5.1 Two tasks scenarios
As expected, using only distillation, as in LwF, does not guarantee good performances for YOLOv8, as it performs even worse than fine-tuning. Indeed, as discussed in Sec. 2, the noisy regression outputs of the teacher prevent the model from learning the new classes and avoiding forgetting. Therefore, a simple L2 loss between the teacher and the student outputs is highly incompatible with the YOLO training.
The results for the Replay method show strong performance on the VOC dataset, particularly in the 10p10 and 19p1 scenarios, where it achieves mAP scores of 54.5 and 60.9, respectively. This indicates that Replay can effectively retain knowledge from earlier tasks when applied to simpler datasets like VOC. When comparing the Replay method with OCDM, we can observe similar results except for the scenario COCO 40p40, which has a slight improvement.
Notably, it performs better than Replay in the 15p1 scenario with a mAP of 50.2, highlighting its effectiveness in handling challenging incremental tasks.
Regarding Pseudo-label, we can observe its effectiveness for most of the scenarios except for the 19p1, where the gap with respect to RCLPOD is significant. We suppose this is due to the low number of old-class objects in the images for the new tasks; the lower the number of classes for the second task, the less overlap between tasks we have (see the plots of Fig. 9 in the Appendix). On the contrary, this is not the case for the 40p40 scenario, where Pseudo-label, despite its simplicity, reaches the best mAP. Even if the overall mAP is higher than the RCLPOD one, a higher forgetting is observed, while plasticity brings that higher final result. For more details, we refer to Sec. .2 Appendix, where we compare RCPLOD and Pseudo-label, the two best methods in the 40p40 scenario, in the case of the bigger model YOLOv8m.
(a) Scenario VOC 15p1 with the mAP50
@@ 5.2 Scenarios VOC 15p1 and COCO 40p10
Here, we discussed the two longer and more challenging streams: VOC 15p1 and COCO 40p10. As mentioned earlier, COCO 40p10 is more complex since the VOC dataset is much smaller. However, the analysis of VOC 15p1 is still meaningful because the overlap in images among tasks is much smaller in VOC than in COCO.
The first observation is that LwF doesn’t perform well with the YOLOv8 architecture for COCO and VOC, achieving comparable performance at Fine-Tuning (lower bound). This confirms the results obtained for the simpler and shorter scenarios.
In contrast to these approaches, Pseudo-Label works very well in a complex scenario like COCO 40p10. This is partially due to the very high intersection between tasks in terms of images, which allows the model to maintain knowledge of previous classes as it implicitly revisits them in the new task. However, in cases where the intersection between tasks is very low the Pseudo-Label performance is expected to be much worse. This is demonstrated by the VOC 15p1 scenario, where performance is low as only one new class is shown, and the intersection with previous tasks is much lower compared to COCO 40p10, where ten new classes are added each time.
Therefore, in practice, we have this interesting duality aspect where Replay for VOC 15p1 performs better than Pseudo-Label and vice versa for the COCO 40p10 scenario.
Unlike Pseudo-label, interference between tasks in Replay is a problem since it implies saving the same image in memory several times with different labels, causing several issues, including an under-exploitation of memory.
OCDM, despite having the same problem of missing annotations, is rewarded by carefully selecting the samples to save, which improves the final performance from 8 for Replay to 11 for OCDM in the COCO 40p10.
Then RCLPOD, which exploits OCDM’s selection mechanism and the Label Propagation technique, receives a significant boost, achieving 20 in the COCO 40p10. This proves how much help the correct management of the information of the samples stored in the replay memory can be effective.
@@ 5.3 Ablation study
Table 2: Ablation study performed to study the effect of each component in RCLPOD, tested on COCO 40p10. LP is the Label Propagation mechanism, OCDM indicates the selection mechanism, FD represents the Feature Distillation component, and Mask represents the Masking Loss component.
Here, we discuss the contribution of each component to the final performance of our novel approach, RCLPOD. Recall that the components are OCDM, Label Propagation (LP), Feature Distillation (FD), and Masking Loss (Mask). Tab. 2 reports the ablation studied, where the effect of each component is studied.
First, we compare Replay (Model 1) with OCDM (Model 2). Even in a short CL scenario like COCO 40p40, OCDM performs better than Replay, showing the positive effect of balancing the Replay memory. By adding the Label Propagation (LP) mechanism (Model 3), a significant improvement of 3 mAP is obtained. In fact, by doing so, we both reduce the interference problem, and we fully utilize the potential of the replay memory. Feature Distillation (Model 4) brings an additional improvement of 0.4 mAP. Once we add the Masking Loss component (Model 5), namely, we use all the components of RCLPOD, we get a further increase of 4.1 mAP. In fact, as shown in Fig.10, the important overlap of many classes between tasks brings interference. Therefore, when a Replay memory is employed, masking is needed even in the case of modern versions of YOLO. Finally, we test OCDM once all the other components are used. By removing OCDM (Model 6), a 1.6 mAP drop is observed, showing the importance of balancing the replay memory in reducing forgetting, particularly for longer streams such as VOC 15p1 and COCO 40p10.
@@ 6 Conclusions and Future Work
To solve the Continual Object Detection, we propose a novel approach called RCLPOD, composed of multiple components such as the Label Propagation technique to exploit, at best, the replay memory, the Feature Distillation to maintain better the old representations, and using a selection mechanism fitted than the classic random mechanism of Replay. By effectively composing these components for the YOLOv8 architecture, we can obtain excellent results for the CLOD setting. We used well-known datasets, such as VOC and COCO, to prove the superiority of RCLPOD to the other methods tested in CLOD.
While our method has allowed us to reduce the gap in the literature with upper-bound Joint Training, it is clear that further work is required to achieve performance comparable to Joint training. An issue that our current approach does not address is that while the introduction of pseudo-labels in replay memory has proven to be beneficial, the noise of such labels must be taken into account as well. Indeed, it is realistic to expect the model to generate false positives and negatives, which will be kept in the replay memory, partially causing noise to the model.
Moreover, we believe that future research should give more attention to longer streams like COCO40p10 and beyond to better represent the real performances of the CL techniques in the CLOD setting.
Appendix
.1 Further training hyper-parameters details
@@
.2 Results with YOLOv8m
@@
.3 Label Distribution for each VOC and COCO scenario
.4 Comparison with literature

### v2

@@ 3.4 Improving Stability
However, the YOLOv8 architecture includes a neck in addition to the backbone $h_{\omega}$. Therefore, we compute the distillation loss on the intermediate features of the backbone and the neck. Recalling that the YOLOv8 backbone outputs three feature maps at multiple levels $C_{3},C_{4},C_{5}$ and the neck does the same $P_{3},P_{4},P_{5}$ [34, 35], the overall loss can be computed as:
@@ 4.1 Scenarios and Metrics
As in [18] to evaluate the performance of YOLOv8, we report the mean average precision (mAP) at the end of the training. Specifically, considering a 0.5 IoU threshold (mAP50) for VOC and the mAP weighted across different IoU from 0.5 to 0.95 (mAP50-95) for COCO.
@@ 4.2 YOLO training details
To evaluate how the CL performances are affected by the model size, in our study we consider two different versions of YOLOv8: YOLOv8n with 3.2M parameters and YOLOv8m with 25.9M parameters. We initialize the backbone parameters with the ones pre-trained on Image-Net, available on [14]. In Table 3 the hyper-parameters used for training are reported. In particular, by following [14, 17], we use SGD with Nestorov momentum and weight decay. For both the learning rate and the momentum we have 3 warmup epochs, while just for the learning rate we employ a linear decay scheduling from $10^{-2}$ to $10^{-4}$. In all the experiments we set the number of epochs per task to 100, which is a suitable value for reaching convergence.
@@ 4.3 CL strategies details
We compare the following methods against RCLPOD: Joint training, Fine-tuning, Replay, OCDM, LwF, and Pseudo-Label. In particular, while Replay, LwF, and Pseudo-label were already evaluated in the CLOD setting, this is the first time, to the best of our knowledge, that the OCDM approach is implemented and tested in the CLOD setting. Join Training is considered as an upper bound and corresponds to the model performances when trained on the entire dataset with all the classes. Fine-tuning is used as a lower bound; each task is trained sequentially, and no CL technique is employed to avoid catastrophic forgetting.
Regarding the Replay methods, the memory has a small capacity fixed to around 5% of the entire dataset, namely 800 images for PASCAL VOC and 6,000 for COCO. For OCDM, we use the same memory size as all the other replay-based approaches. Moreover, we provide additional results for the replay-based approaches in Sec. A.5 by changing the memory size to evaluate its impact on performance.
@@
Fine-Tuning 36.4 16.5 16.4 2.9 13.6 2.9
Replay [28] 54.5 60.8 56.9 40.0 14.1 8.0
LwF [20] 57.3 60.0 57.0 41.5 19.0 16.2
OCDM [31] 56.5 59.9 54.7 50.2 16.7 11.0
Pseudo-label [27] 71.4 46.6 59.7 16.3 24.4 16.2
@@ 5 Results
In this section, we discuss the outcomes of each method on the VOC and COCO CLOD benchmarks. Table 1 reports the results after all tasks have been completed for YOLOv8n. Additional results for YOLOv8M are provided in Sec. A.2. In Sec. 5.1, we discuss the results obtained in the 2-tasks scenarios, namely 15p5, 10p10, 19p1, and 40p40. In Sec. 5.2, we consider the more challenging scenarios 15p1 and 40p10.
@@ 5.1 Two tasks scenarios
Replay shows good performance on the VOC dataset, particularly in the 10p10 and 19p1 scenarios, where it achieves mAP scores of 54.5 and 60.9, respectively. This indicates that Replay can effectively retain knowledge from earlier tasks when applied to simpler datasets like VOC. When comparing the Replay method with OCDM, we can observe similar results except for the scenario COCO 40p40, which has a slight improvement.
Notably, OCDM performs better than Replay in the long 15p1 scenario with a mAP of 50.2, highlighting its effectiveness in handling challenging incremental tasks.
Regarding Pseudo-label, we can observe its effectiveness for most of the scenarios except for the 19p1, where the gap to RCLPOD is significant. We suppose this is due to the low number of old-class objects in the images for the new tasks; the lower the number of classes for the second task, the less overlap between tasks we have (see the plots of Fig. 9 in the Appendix). On the contrary, this is not the case for the 40p40 scenario, where Pseudo-label, despite its simplicity, reaches the best mAP. Even if the overall mAP is higher than the RCLPOD one, a higher forgetting is observed, while plasticity brings that higher final result. For more details, we refer to Sec. A.2 Appendix, where we compare RCPLOD and Pseudo-label, the two best methods in the 40p40 scenario, in the case of the bigger model YOLOv8m.
For what concerns LwF, the results obtained show that this method can obtain good performance in many scenarios. However, we can observe the model plasticity is significantly reduced. This is due to the noisy regression output of the teacher, then the loss between the teacher and the student output prevents the model from acquiring new knowledge. For a more detailed discussion, we refer to Sec. A.4 in the Appendix.
(a) Scenario VOC 15p1 with the mAP50.
@@ 5.2 Scenarios VOC 15p1 and COCO 40p10
[…]
As highlighted in Tab 1, Pseudo-Label works very well in a complex scenario like COCO 40p10. This is partially due to the very high intersection between tasks in terms of images, which allows the model to maintain knowledge of previous classes as it implicitly revisits them in the new task. However, in cases where the intersection between tasks is very low the Pseudo-Label performance is expected to be much worse. This is demonstrated by the VOC 15p1 scenario, where performance is low as only one new class is shown, and the intersection with previous tasks is much lower compared to COCO 40p10, where ten new classes are added each time.
Therefore, in practice, we have this interesting duality aspect where Replay for VOC 15p1 performs better than Pseudo-Label and vice versa for the COCO 40p10 scenario. Unlike Pseudo-label, interference between tasks in Replay is a problem since it implies saving the same image in memory several times with different labels, causing several issues, including an under-exploitation of memory. OCDM, despite having the same problem of missing annotations, is rewarded by carefully selecting the samples to save, which improves the final performance from 8 for Replay to 11 for OCDM in the COCO 40p10.
Then RCLPOD, which enhances the replay memory by improving the ground truth associated with the samples, receives a significant boost, achieving 20 in the COCO 40p10. This proves how much help the correct management of the information of the samples stored in the replay memory can be effective. Finally, even if LwF reaches good performance in both scenarios, the same problem highlighted for shorter tasks, namely, that the model achieves poor performance for new tasks is observed.
@@ 5.3 Ablation study
Table 2: Ablation study performed to study the effect of each component in RCLPOD, tested on COCO 40p10. LP is the Label Propagation mechanism, OCDM indicates the selection mechanism, FD represents the Feature Distillation component, and Mask represents the masking component.
Here, we discuss the contribution of each part of our RCLPOD approach to the final performance. Specifically, we identify four main parts. The Label Propagation (LP) mechanism described in Sec. 3.1 to enhance the replay memory. The selection mechanism (OCDM) to balance the class distribution of the replay memory that promotes class fairness. The masking (Mask) introduced in Sec. 3.3 to reduce the task interference in the case of overlapping objects for the YOLO architecture. Finally, Feature Distillation (FD) is applied to low-level representations to improve the stability and reduce forgetting described in Sec. 3.4.
Tab. 2 reports the ablation results. First, we compare Replay (Model 1) with OCDM (Model 2). Even in a short CL scenario like COCO 40p40, OCDM performs better than Replay, showing the positive effect of balancing the Replay memory. By adding the Label Propagation (LP) mechanism (Model 3), a significant improvement of 3 mAP is obtained. In fact, by doing so, we both reduce the interference problem, and we fully utilize the potential of the replay memory. Feature Distillation (Model 4) brings an additional improvement of 0.4 mAP. Once we add the masking component (Model 5), using all the components of RCLPOD, we get a further increase of 4 mAP. In fact, as shown in Fig.10, the important overlap of many classes between tasks brings interference. Therefore, when a Replay memory is employed, masking is needed even in the case of modern versions of YOLO. Finally, we test OCDM once all the other components are used. By removing OCDM (Model 6), a 1.4 mAP drop is observed, showing the importance of balancing the replay memory in reducing forgetting, particularly for longer streams such as VOC 15p1 and COCO 40p10.
@@ 6 Conclusions and Future Work
To improve the performances of Continual Object Detection, we propose a novel approach called RCPLOD. Our method, enhances the replay memory through class distribution balancing and Label Propagation, using existing data more effectively while solving the missing annotations and the task interference problem. RCLPOD outperforms existing techniques on well-established benchmarks such as VOC and COC.
The new enhanced memory strategy described in this work provides an alternative to distillation-based approaches and can be reapplied to other architectures. In addition, our algorithm doesn’t occupy more memory than other approaches like Replay and LwF. Similarly, while the computing could be slightly higher than Replay due to the procedure to enhance the replay memory, it is still a lightweight approach since it just needs to modify the information stored in the replay memory. Moreover, our approach is developed to work with modern architectures like YOLOv8, making it suitable for dynamic, real-world applications such as autonomous driving [4] and robotics [3], where continuous learning and resource efficiency are essential.
As future research, we believe that our approach can be used as a base for the proposal of new advancements in replay-based approaches. Indeed, the Label Propagation mechanism tested in the CLOD scenario proves its efficacy, making it a suitable base to build upon, since it facilitates more stable and memory-efficient models. In particular, the base of our strategy is designed to work independently of specific detector architectures. The technique manages memory and labeling in ways that are generally compatible with any object detection model using replay-based continual learning, simplifying the extension to other architectures. In addition, future research should give more attention to longer and more complex streams like COCO40p10 and beyond to better represent the real performances of the CL techniques in the CLOD setting. Following this direction, streams closer to realistic scenarios should be considered to represent better the challenges of Continual Object Detection.
Appendix A Appendix
A.1 Further training hyper-parameters details
@@
A.2 Results with YOLOv8m
A.3 Label Distribution for each VOC and COCO scenario
A.4 LwF and stability-plasticity dilemma
As shown in Tab. 5, LwF is effective for reducing forgetting but prevents the model from learning a new task. Moreover, in contrast to Pseudo-Label, it prevents forgetting even in the VOC19p1 scenario, where, as shown in the previous section, objects for previous tasks may not appear in the dataset of the new task. These two results are justified by the noisy output of the teacher: contrary to inference, all those predictions for which the model is highly uncertain (very low classification score for any class) are used as target for the student. Therefore, independently by the choice of $\lambda$, LwF constraints the student to mimic the teacher regression output even when it is not necessary.
A.5 Memory efficiency
In this section, we provide additional results for the replay-based approaches by changing the memory size to evaluate their impact on performance. In particular, we compare the two best replay-based methods, RCLPOD and OCDM, in the VOC15p1 scenario, where replay-based methods are effective. From Tab. 6, we can observe that by decreasing the memory size $m$ from $m=800$ (5%) to $m=400$ (2.5%), RCLPOD is less sensitive to this hyper-parameter than OCDM, namely the performance decay slower w.r.t. a memory size decrease.
A.6 Comparison with literature
A.7 Optimizing Class Distribution in Memory (OCDM)
Optimizing Class Distribution in Memory (OCDM) is a greedy approach that selects a subset of samples such that the final distribution of the labels in memory is as close as possible to a uniform target distribution. OCDM formulates the memory update mechanism as an optimization problem. This greedy algorithm is detailed in Alg. 1.
In particular, once a new batch of data $\mathcal{B}_{t}$ of size $b_{t}$ for task $t\in\mathcal{T}$ arrives, the algorithm updates the memory $\mathcal{M}$ of size $M$ by solving the following optimization problem:
$\min_{\Omega}$ $d(\mathbb{p},\mathbb{p}_{\Omega})$ (5)
subject to $\Omega\subseteq\mathcal{M}\cup\mathcal{B}_{t}$
$|\Omega|=M$
where p represents the target distribution i.e. the ideal optimal solution, while, $\textbf{p}_{\Omega}$ represents the distribution of the labels produced from the samples of the set $\Omega$. The $d(\cdot,\cdot)$ function used to measure the difference between the two distributions is the Kullback–Leibler (KL) divergence. The target distribution proposed in [39] is defined as follows:
$p_{i}=\frac{(n_{i})^{\rho}}{\sum_{j=1}^{C}(n_{j})^{\rho}}$ (6)
where $n_{i}$ is the frequency of a class $i$ and $\rho$ is the allocation power. Using $\rho=0$ the samples are saved in memory $\mathcal{M}$ in order to have equally distributed classes.
@@ A.7 Optimizing Class Distribution in Memory (OCDM)
CL Scenario LwF RCLPOD Pseudo-Label
old new all old new all old new all
VOC10p10 50.7 63.8 57.3 70.9 74.1 72.5 68.5 74.3 71.4
VOC19p1 60.6 50.3 60.0 68.8 61.8 68.4 45.8 61.9 46.6
VOC15p5 61.5 43.4 57.0 71.6 55.8 67.7 61.1 55.7 59.7
COCO40p40 26.9 11.0 19.0 25.7 22.5 24.1 22.7 26.0 24.3
Table 5: Comparison LwF, RCLPOD and Pseudo-Label for 2-tasks scenarios. "old" is the mAP for the classes of Task 1 after Task 2 training, "new" is the mAP for the new classes, while "all" is the mAP over all the classes.
CL Method $m=800$ $m=400$ $\Delta\%\;(\downarrow)$
RCLPOD 58.3 52.5 10.0%
OCDM 50.2 43.4 13.5%
Table 6: Results for the best replay-based methods in the VOC15p1 scenario varying the memory size $m$.
@@
Input : task stream $\mathcal{T}$, total size of memory M
${}\mathcal{M}\leftarrow\{\}$
  /* Initialize the memory */
for $t\in\mathcal{T}$ do
       $\mathcal{D}_{t}\leftarrow$ Get Dataset $\mathcal{D}_{t}=\{X_{t},Y_{t}\}$ of task $t$
      for $B_{t}\in\mathcal{D}_{t}$ do
             if $|\mathcal{M}|\leq M$ then
                   $\text{diff}\leftarrow|\mathcal{M}|-M$
                   $V_{t}\leftarrow$ select randomly $\min(|B_{t}|,\text{diff})$ samples from $B_{t}$
                   $B_{t}\leftarrow B_{t}\setminus V_{t}$
                   $\mathcal{M}\leftarrow\mathcal{M}\cup V_{t}$
             end if
            if $|B_{t}|>0$ then
                   $\Omega\leftarrow\mathcal{M}\cup B_{t}$
                   for $k\in[1,2,...,|B_{t}|]$ do
                         $i=arg\min_{j\in\Omega}d(\mathbb{p},\mathbb{p}_{\Omega\setminus\{j\}})$
                        $\Omega\leftarrow\Omega\setminus\{i\}$
                   end for
                  $\mathcal{M}\leftarrow\Omega$
             end if
       end for
end for
Algorithm 1 Optimizing Class Distribution in Memory (OCDM)